Blog Archive

Last modified by Michael Sibayan on 2015/02/13 12:22
Comments (0) · Attachments (0) · History

Blog posts for October 2018

10-4-2018 LEO Tipping bucket recalibration

We have recalibrated the LEO Novalynx tipping bucket. 10/4/2018 10:45pm data set onwards are using the new calibration values.

novalynx_calibration.png

6/4/2019: The following shows the updated calibration coefficients used on the Novalynx.

Novalynx_recalibrations.PNG

PE102-038 East Bay -2 Swap

The PE102-038 in East bay -2 position was measuring about 1/2 of expected counts. Flowmeter from the +4 position in East bay was swapped with -2 position until a replacement is available. 

Tags:
Created by Michael Sibayan on 2015/02/13 12:22

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation